OpenStudy (anonymous):

Cylinder A has a radius of 12 inches and a height of 6 inches. Cylinder B has a volume of 648π. What is the percent change in volume between cylinders A and B? Cylinder B is 50% smaller than cylinder A. Cylinder B is 25% smaller than cylinder A. Cylinder B is 150% bigger than cylinder A. Cylinder B is 200% bigger than cylinder A.

OpenStudy (anonymous):

find the volume of cylinder A

OpenStudy (anonymous):

3.14 * 12^2 *6 3.14 * 144 *6 2712.96

OpenStudy (anonymous):

yup, now find the volume of cylinder B

OpenStudy (anonymous):

then divide the bigger volume by the smaller volume

OpenStudy (anonymous):

then convert that decimal into a percent

OpenStudy (anonymous):

but how do i find the volume of cylinder B when i dont know the radius

OpenStudy (anonymous):

because it says the volume of cylinder B is 648 times pi

OpenStudy (anonymous):

648 * 3.14 + 2034.72

OpenStudy (anonymous):

no,volume of A\[=\pi 12^2*6=864 \pi\] volume of B\[=648 \pi\] change in volume\[=864 \pi-648 \pi=216 \pi\] % change\[=\frac{ 216 \pi }{ 864 \pi }\times 100=25 % \] =25 %

OpenStudy (anonymous):

how did you get the volume of A to be 864.. Im confused ? @surjithayer

OpenStudy (anonymous):

\[v= \pi r^2 h\]

OpenStudy (anonymous):

he/she didn't get 864, he/she got 864*pi

OpenStudy (anonymous):

Oh i see what you did now my bad(:

OpenStudy (anonymous):

do not substitute the value of pi as the volume of B is 648 pi.

OpenStudy (anonymous):

oh, never mind - listen to @surjithayer I might be wrong

OpenStudy (anonymous):

i am a man.

OpenStudy (anonymous):

I understand now, Thank you soooooo muchhh!!

OpenStudy (anonymous):

yw